# DSSmoothing: Toward Certified Dataset Ownership Verification for Pre-trained Language Models via Dual-Space Smoothing

### v1

@@ Abstract.
Large web-scale datasets have driven the rapid advancement of pre-trained language models (PLMs), but unauthorized data usage has raised serious copyright concerns. Existing dataset ownership verification (DOV) methods typically assume that watermarks remain stable during inference; however, this assumption often fails under natural noise and adversary-crafted perturbations. We propose the first certified dataset ownership verification method for PLMs based on dual-space smoothing ($i.e.$, DSSmoothing). To address the challenges of text discreteness and semantic sensitivity, DSSmoothing introduces continuous perturbations in the embedding space to capture semantic robustness and applies controlled token reordering in the permutation space to capture sequential robustness. DSSmoothing consists of two stages: in the first stage, triggers are collaboratively embedded in both spaces to generate norm-constrained and robust watermarked datasets; in the second stage, randomized smoothing is applied in both spaces during verification to compute the watermark robustness (WR) of suspicious models and statistically compare it with the principal probability (PP) values of a set of benign models. Theoretically, DSSmoothing provides provable robustness guarantees for dataset ownership verification by ensuring that WR consistently exceeds PP under bounded dual-space perturbations. Extensive experiments on multiple representative web datasets demonstrate that DSSmoothing achieves stable and reliable verification performance and exhibits robustness against potential adaptive attacks.
Certified Dataset Watermarking, Dataset Ownership Verification, Pre-trained Language Models, AI Copyright Protection, Data Protection
copyright: noneccs: Security and privacy Data security and privacyccs: Security and privacy Data transparency and provenance
@@ 1. Introduction
To the best of our knowledge, dataset ownership verification (DOV) (Li et al., 2023; Guo et al., 2023; Wei et al., 2024; Guo et al., 2025; Qiao et al., 2025) serves as an effective post-hoc auditing (Du et al., 2025; Li et al., 2025a) method that determines whether a suspicious model has been trained on a protected dataset by examining its prediction behaviors on specially designed verification samples under black-box settings. The process generally involves two stages: dataset watermarking and ownership verification. In the first stage, dataset owners make small, deliberate modifications to a subset of samples to generate a watermarked version that preserves normal accuracy on benign tests while inducing distinctive, predefined behaviors on verification samples. In the second stage, defenders ($i.e.$, dataset owners) query the suspicious model through API access and check whether such behaviors occur. If detected, it can be concluded that the model was indeed trained on the protected dataset.
@@
The answer to this question is affirmative, although certified watermarking methods from the image domain cannot be directly transferred to pre-trained language models (PLMs). The difficulty mainly lies in two aspects. First, this method relies on well-defined continuous perturbations to enable noise-based certification, whereas text consists of discrete tokens that cannot accommodate smooth Gaussian noise as pixels do. Second, small perturbations in images typically preserve visual content, while text is highly semantically sensitive, and even minor modifications may disrupt linguistic fluency or alter meaning. To address these challenges, we propose DSSmoothing, the first certified dataset watermarking method for PLMs based on dual-space smoothing (as illustrated in Figure 1). In general, our DSSmoothing operates in two stages. In the first stage, inspired by TextCRS (Zhang et al., 2024), it collaboratively embeds triggers in both the embedding and permutation spaces to generate norm-constrained and robust watermarked datasets. This design jointly addresses the aforementioned challenges: in the embedding space, continuous perturbations are applied to map discrete tokens into continuous vector space, thereby capturing semantic robustness and re-enabling noise-based certification; in the permutation space, controlled token reordering is performed to enhance sequential robustness while maintaining linguistic fluency and semantic consistency. In the second stage, DSSmoothing applies smoothing perturbations in both spaces to obtain the watermark robustness (WR) of the suspicious model and the principal probabilities (PP) of a set of benign models. The computed WR and PP metrics support our theoretical analysis, which proves that under bounded dual-space perturbations, WR (defined on watermarked samples) remains strictly higher than PP (defined on clean samples). Therefore, when the WR of a suspicious model significantly exceeds the majority of PP values from benign models, we can reliably conclude that the model has been trained on the protected dataset.
@@ 2.1. Dataset Auditing
Dataset auditing (DA) (Huang et al., 2024; Du et al., 2025) aims to determine whether a dataset has been used to train a suspicious model, and can be categorized into non-intrusive and intrusive approaches. Non-intrusive methods leave the dataset unchanged but often suffer from high false-positive rates (Shao et al., 2025; Li et al., 2025d). Therefore, this paper focuses on intrusive approaches, particularly dataset ownership verification (DOV) (Li et al., 2022a; Guo et al., 2024). DOV introduces imperceptible watermark samples into the original dataset to construct a watermarked version, enabling ownership verification in a black-box setting. These samples preserve model performance on clean test data while inducing distinctive behaviors on verification inputs. Research has evolved from poisoned-label backdoors (Li et al., 2023) to clean-label backdoors (Tang et al., 2023), and further to harmless watermarks (Guo et al., 2023), with extensions to point clouds (Wei et al., 2024), text-to-image diffusion models (Li et al., 2025b), and speech data (Li et al., 2025c; Guo et al., 2025). More recently, attention has shifted to pre-trained language models (PLMs) (Liu et al., 2023), where textual backdoors remain the dominant strategy (see Appendix B.1). However, existing DOV techniques lack theoretical robustness guarantees and are vulnerable to adaptive attacks. Recently, Qiao et al. (Qiao et al., 2025) proposed the first certified dataset watermarking (CertDW) method, achieving provable robustness under bounded perturbations in the image domain, as detailed in Section 2.2. Yet, certified dataset ownership verification for textual data remains challenging due to discreteness and semantic sensitivity. Motivated by this gap, this paper aims to develop a certified DOV method for textual data.
2.2. Random Smoothing and CertDW
Randomized Smoothing (RS) (Cohen et al., 2019) was originally introduced for adversarial defense in image classification. Its core idea is to add Gaussian noise to the input multiple times and aggregate predictions via majority voting, ensuring stable outputs within a bounded perturbation (see Appendix B.2). Later studies (Bansal et al., 2022; Jiang et al., 2023; Ren et al., 2023) extended RS to model watermarking by injecting noise into the parameter space, but these methods rely on training details and remain limited to model-level copyright protection. Building on this line of work, Qiao et al. proposed CertDW (Qiao et al., 2025), the first to apply RS to dataset ownership verification. CertDW provides robustness guarantees in the image sample space, showing that ownership verification remains reliable under malicious attacks as long as certain conditions are satisfied, as detailed below.
@@ 2.2. Random Smoothing and CertDW
[…]
@@ 4.1. Preliminaries
In this subsection, we first review the main pipeline of text dataset ownership verification (DOV) methods, and then introduce the key definition and lemma that form the theoretical foundation of our proposed DSSmoothing method.
Main Pipeline of Text Classification DOV Methods. Consider a standard text classification dataset $\mathcal{D}=\{(\bm{x}_{n},y_{n})\}_{n=1}^{N}$ across $K$ classes, where each text sequence $\bm{x}_{n}\in\mathcal{X}$ is associated with a label $y_{n}\in\mathcal{Y}=\{1,2,\ldots,K\}$. During the watermarking phase, the dataset owner constructs a watermarked dataset $\mathcal{D}_{w}$ from $\mathcal{D}$. Specifically, $\mathcal{D}_{w}=\mathcal{D}_{m}\cup\mathcal{D}_{r}$, where $\mathcal{D}_{m}$ contains modified samples from a small selected subset $\mathcal{D}_{s}\subset\mathcal{D}$, and $\mathcal{D}_{r}=\mathcal{D}\setminus\mathcal{D}_{s}$ represents the remaining benign samples. The modified subset $\mathcal{D}_{m}$ is generated through a dataset-specific text generator $G_{X}:\mathcal{X}\rightarrow\mathcal{X}$ and a label generator $G_{Y}:\mathcal{Y}\rightarrow\mathcal{Y}$, defined as $\mathcal{D}_{m}=\{(\hat{\bm{x}},\hat{y}):\hat{\bm{x}}=G_{X}(\bm{x}),\hat{y}=G_{Y}(y),(\bm{x},y)\in\mathcal{D}_{s}\}$. For instance, in BadWord-based DOV (Li et al., 2023), $G_{X}$ inserts specific trigger words ($e.g.$, “cf”, “mn”) into the text, and $G_{Y}(y)=\hat{y}$ assigns a fixed target label. In AddSent-based DOV (Dai et al., 2019), $G_{X}$ appends a trigger sentence ($e.g.$, “I watch this 3D movie”) to the original text. The watermarking rate is defined as $\gamma\triangleq\frac{|\mathcal{D}_{m}|}{|\mathcal{D}_{w}|}$. During the ownership verification phase, the dataset owner verifies whether a suspicious model $f(\cdot;\bm{\theta}):\mathcal{X}\rightarrow\mathcal{Y}$ has been trained on the protected dataset $\mathcal{D}_{w}$ by querying it with specially designed verification samples in a black-box setting.
@@
4.2. Threat Model.
In this paper, we adopt the classical DOV setting, where the dataset owner embeds trigger-based watermarks into the text before release, and downstream users may train PLMs on it. During the verification phase, the owner can only query the suspicious model in a black-box setting. Unlike prior work that implicitly assumes watermark embeddings remain stable during verification, we explicitly consider two realistic perturbations: stochastic noise that naturally arises during inference, and worst-case perturbations applied to word embeddings to evade ownership verification while preserving semantics. This threat model better reflects web-scale deployment environments, thereby providing a more reliable basis for evaluating verification effectiveness.
@@ 4.5. Ownership Verification of DSSmoothing
(6) $\small p_{k}\!\big((\bm{U},\bm{W})\mid g_{w},\mathcal{P}_{\rho,\varepsilon}\big)\approx\frac{1}{M}\sum_{i=1}^{M}\mathbb{I}\!\Big\{\arg\max g\big(\psi(\bm{U},\rho_{i})\cdot\phi(\bm{W},\varepsilon_{i});w\big)=k\Big\},$
@@ 5.1. Main Settings
Dataset and Model Selections. We evaluate our method on two widely-used text classification datasets that exemplify web-collected data: AG’s News (AG) (Zhang et al., 2015), a news classification dataset aggregated from over 2,000 online news sources with four topic categories (World, Sports, Business, and Sci/Tech), representing professionally curated journalism content; and SST-2 (Socher et al., 2013), a binary sentiment classification dataset derived from movie reviews scraped from Rotten Tomatoes, exemplifying user-generated web content. To validate the generalizability across different architectures, we conduct experiments on three representative pre-trained language models: BERT-base-uncased (Devlin et al., 2019), GPT-2 base (Radford et al., 2019), and OPT-1.3B (Zhang et al., 2022).
@@
Figure 10. WSR of our method’s watermarked samples under perturbations in the embedding space. $d_{N}$ and $d_{A}$ denote the noise and adversary-crafted directions, respectively, and $\bullet$’ marks the original watermarked sample. Left: BadWord; right: AddSent.
@@ 6. Conclusion
In this paper, we revisited existing dataset ownership verification (DOV) methods for pre-trained language models (PLMs) and revealed their vulnerabilities under natural noise and deliberately adversary-crafted perturbations. To address these issues, we proposed DSSmoothing, the first certified dataset watermarking method for PLMs based on dual-space smoothing. By introducing controlled perturbations in both embedding and permutation spaces, DSSmoothing captures semantic and sequential robustness simultaneously. Our theoretical analysis proves that under bounded perturbations, a strict lower bound exists between watermark robustness (WR) and principal probability (PP), thereby enabling provable ownership verification. Extensive experiments on benchmark datasets validate the effectiveness of DSSmoothing, showing that it not only outperforms existing methods but also resists potential adaptive attacks. These findings highlight DSSmoothing as a novel perspective for copyright protection of large-scale text datasets, offering a feasible and theoretically guaranteed solution for trustworthy dataset sharing and trading.

### v2

@@ Abstract.
Large web-scale datasets have driven the rapid advancement of pre-trained language models (PLMs), but unauthorized data usage has raised serious copyright concerns. Existing dataset ownership verification (DOV) methods typically assume that watermarks remain stable during inference; however, this assumption often fails under natural noise and adversary-crafted perturbations. We propose the first certified dataset ownership verification method for PLMs under a gray-box setting ($i.e.$, the defender can only query the suspicious model but is aware of its input representation module), based on dual-space smoothing ($i.e.$, DSSmoothing). To address the challenges of text discreteness and semantic sensitivity, DSSmoothing introduces continuous perturbations in the embedding space to capture semantic robustness and applies controlled token reordering in the permutation space to capture sequential robustness. DSSmoothing consists of two stages: in the first stage, triggers are collaboratively embedded in both spaces to generate norm-constrained and robust watermarked datasets; in the second stage, randomized smoothing is applied in both spaces during verification to compute the watermark robustness (WR) of suspicious models and statistically compare it with the principal probability (PP) values of a set of benign models. Theoretically, DSSmoothing provides provable robustness guarantees for dataset ownership verification by ensuring that WR consistently exceeds PP under bounded dual-space perturbations. Extensive experiments on multiple representative web datasets demonstrate that DSSmoothing achieves stable and reliable verification performance and exhibits robustness against potential adaptive attacks. Our code is available at https://github.com/NcepuQiaoTing/DSSmoothing.
Certified Dataset Watermark, Dataset Ownership Verification, Pre-trained Language Models, AI Copyright Protection, Data Protection
journalyear: 2026copyright: ccconference: Proceedings of the ACM Web Conference 2026; April 13–17, 2026; Dubai, United Arab Emiratesbooktitle: Proceedings of the ACM Web Conference 2026 (WWW ’26), April 13–17, 2026, Dubai, United Arab Emiratesdoi: 10.1145/3774904.3792457isbn: 979-8-4007-2307-0/2026/04ccs: Security and privacy Digital rights managementccs: Security and privacy Privacy-preserving protocols
@@ 1. Introduction
Conceptual comparison figure contrasting existing dataset ownership verification methods for pre-trained language models with DSSmoothing. The figure highlights that DSSmoothing uses dual-space smoothing in embedding and permutation spaces to improve robustness against natural noise and adversarial perturbations.
@@
To the best of our knowledge, dataset ownership verification (DOV) (Li et al., 2023; Guo et al., 2023; Wei et al., 2024; Guo et al., 2025; Qiao et al., 2025) serves as an effective post-hoc auditing (Du et al., 2025; Li et al., 2025b) method that determines whether a suspicious model has been trained on a protected dataset by examining its prediction behaviors on specially designed verification samples. The process generally involves two stages: dataset watermarking and ownership verification. In the first stage, dataset owners make small, deliberate modifications to a subset of samples to generate a watermarked version that preserves normal accuracy on benign tests while inducing distinctive, predefined behaviors on verification samples. In the second stage, defenders ($i.e.$, dataset owners) query the suspicious model through API access and check whether such behaviors occur. If detected, it can be concluded that the model was indeed trained on the protected dataset.
@@
The answer to this question is affirmative, although certified watermarking methods from the image domain cannot be directly transferred to PLMs. The difficulty mainly lies in two aspects. First, this method relies on well-defined continuous perturbations to enable noise-based certification, whereas text consists of discrete tokens that cannot accommodate smooth Gaussian noise as pixels do. Second, small perturbations in images typically preserve visual content, while text is highly semantically sensitive, and even minor modifications may disrupt linguistic fluency or alter meaning. To address these challenges, we propose DSSmoothing, the first certified dataset watermarking method for PLMs based on dual-space smoothing (as illustrated in Figure 1). The method is designed under a gray-box setting, where the defender has the knowledge of its input representation module but can only query the suspicious model. In general, our DSSmoothing operates in two stages. In the first stage, inspired by TextCRS (Zhang et al., 2024), it collaboratively embeds triggers in both the embedding and permutation spaces to generate norm-constrained and robust watermarked datasets. This design jointly addresses the aforementioned challenges: in the embedding space, continuous perturbations are applied to map discrete tokens into continuous vector space, thereby capturing semantic robustness and re-enabling noise-based certification; in the permutation space, controlled token reordering is performed to enhance sequential robustness while maintaining linguistic fluency and semantic consistency. In the second stage, DSSmoothing applies smoothing perturbations in both spaces to obtain the watermark robustness (WR) of the suspicious model and the principal probabilities (PP) of a set of benign models. The computed WR and PP metrics support our theoretical analysis, which proves that under bounded dual-space perturbations, WR (defined on watermarked samples) remains strictly higher than PP (defined on clean samples). Accordingly, when the WR of a suspicious model significantly exceeds the majority of benign-model PP values, we can conclude that the model has been trained on the protected dataset.
@@ 2.1. Dataset Auditing
Dataset auditing (DA) (Huang et al., 2024; Du et al., 2025) aims to determine whether a dataset has been used to train a suspicious model, and can be categorized into non-intrusive and intrusive approaches. Non-intrusive methods leave the dataset unchanged but often suffer from high false-positive rates (Shao et al., 2025; Li et al., 2025d). Therefore, this paper focuses on intrusive approaches, particularly dataset ownership verification (DOV) (Li et al., 2022a; Guo et al., 2024). DOV introduces imperceptible watermark samples into the original dataset to construct a watermarked version, enabling ownership verification typically under black-box access ($e.g.$, via API queries). These samples preserve model performance on clean test data while inducing distinctive behaviors on verification inputs. Research has evolved from poisoned-label backdoors (Li et al., 2023) to clean-label backdoors (Tang et al., 2023), and further to harmless watermarks (Guo et al., 2023), with extensions to point clouds (Wei et al., 2024), text-to-image diffusion models (Li et al., 2025a), and speech data (Li et al., 2025c; Guo et al., 2025). More recently, attention has shifted to pre-trained language models (PLMs) (Liu et al., 2023), where textual backdoors remain the dominant strategy (see Appendix B.1). However, existing DOV methods generally lack provable robustness guarantees and are vulnerable to adaptive attacks. While CertDW (Qiao et al., 2025), a certified dataset watermarking approach, represents an important first step by establishing provable robustness guarantees under bounded perturbations in the image domain, extending such certification to textual data remains fundamentally challenging due to the discreteness of text and its sensitivity to semantic constraints.
@@
Two subfigures comparing watermark robustness (WR) under embedding-space perturbations for two trigger types. The left panel corresponds to BadWord and the right panel to AddSent; in each panel, the original watermarked sample is marked with a bullet and the curves show WR under noise and adversary-crafted directions.
2.2. Random Smoothing and CertDW
Randomized Smoothing (RS) (Cohen et al., 2019) was originally proposed for adversarial defense in image classification. By repeatedly injecting Gaussian noise into the input and aggregating predictions via majority voting, RS guarantees stable outputs within a bounded perturbation radius (see Appendix B.2). Later studies (Bansal et al., 2022; Jiang et al., 2023; Ren et al., 2023) extended RS to model watermarking by injecting noise into the parameter space to protect model ownership, but these methods heavily rely on training details and are largely limited to the parameter level. Building on this line of research, Qiao et al. proposed CertDW (Qiao et al., 2025), the first approach to apply RS to dataset ownership verification, providing robustness guarantees in the image sample space. Under certain conditions, ownership verification remains reliable even in the presence of malicious attacks (see Appendix B.3). Nevertheless, this formalization is difficult to directly extend to text data: discrete tokens cannot naturally accommodate continuous Gaussian noise, rendering $\ell_{p}$-norm-based perturbation bounds inapplicable, and textual perturbations must preserve both linguistic fluency and semantic coherence. These fundamental challenges call for a new certification mechanism tailored to the discrete and semantically sensitive nature of textual data, which is the focus of this work.
@@ 3. Revisiting PLM Dataset Watermarking
Two subfigures comparing watermark robustness (WR) under noise for two trigger types: BadWord on the left and AddSent on the right.
@@ 3.2. Vulnerability under Adversary-Crafted Perturbations
Overview of the DSSmoothing pipeline. The figure illustrates the two-stage process: watermark embedding in dual spaces to generate a robust watermarked dataset, and dual-space randomized smoothing during verification to compute watermark robustness for a suspicious model and compare it against the principal probability of benign models.
@@ 4.1. Preliminaries
Main Pipeline of Text Classification DOV Methods. Consider a standard text classification dataset $\mathcal{D}=\{(\bm{x}_{n},y_{n})\}_{n=1}^{N}$ across $K$ classes, where each text sequence $\bm{x}_{n}\in\mathcal{X}$ is associated with a label $y_{n}\in\mathcal{Y}=\{1,2,\ldots,K\}$. During the watermarking phase, the dataset owner constructs a watermarked dataset $\mathcal{D}_{w}$ from $\mathcal{D}$. Specifically, $\mathcal{D}_{w}=\mathcal{D}_{m}\cup\mathcal{D}_{r}$, where $\mathcal{D}_{m}$ contains modified samples from a small selected subset $\mathcal{D}_{s}\subset\mathcal{D}$, and $\mathcal{D}_{r}=\mathcal{D}\setminus\mathcal{D}_{s}$ represents the remaining benign samples. The modified subset $\mathcal{D}_{m}$ is generated through a dataset-specific text generator $G_{X}:\mathcal{X}\rightarrow\mathcal{X}$ and a label generator $G_{Y}:\mathcal{Y}\rightarrow\mathcal{Y}$, defined as $\mathcal{D}_{m}=\{(\hat{\bm{x}},\hat{y}):\hat{\bm{x}}=G_{X}(\bm{x}),\hat{y}=G_{Y}(y),(\bm{x},y)\in\mathcal{D}_{s}\}$. For instance, in BadWord-based DOV (Li et al., 2023), $G_{X}$ inserts specific trigger words ($e.g.$, “cf”, “mn”) into the text, and $G_{Y}(y)=\hat{y}$ assigns a fixed target label. In AddSent-based DOV (Dai et al., 2019), $G_{X}$ appends a trigger sentence ($e.g.$, “I watch this 3D movie”) to the original text. The watermarking rate is defined as $\gamma\triangleq\frac{|\mathcal{D}_{m}|}{|\mathcal{D}_{w}|}$. During the ownership verification phase, the dataset owner verifies whether a suspicious model $f(\cdot;\bm{\theta}):\mathcal{X}\rightarrow\mathcal{Y}$ has been trained on the protected dataset $\mathcal{D}_{w}$ by querying it with specially designed verification samples typically under the black-box setting.
@@
4.2. Threat Model
We consider a standard DOV pipeline for PLMs, where a dataset owner embeds trigger-based watermarks into a text dataset prior to release, and downstream users may subsequently train PLMs on the released data. At verification time, the owner seeks to determine whether a suspicious model has been trained on the protected dataset. We assume a query-only verification setting: defender ($i.e.$, data owner) can interact with the suspicious model via API, but has no access to its parameters or training details, and cannot modify the model. Unlike a strict black-box setting, we assume that the defender is aware of the input representation module adopted by the model, which maps discrete text inputs into continuous embeddings. This assumption is standard for modern PLMs, where tokenization and embedding mappings are publicly available or reproducible.
@@ 4.5. Ownership Verification of DSSmoothing
(6) $p_{k}\!\big((\bm{U},\bm{W})\mid g_{w},\mathcal{P}_{\rho,\varepsilon}\big)\approx\frac{1}{M}\sum_{i=1}^{M}\mathbb{I}\!\Big\{\arg\max g\big(\psi(\bm{U},\rho_{i})\cdot\phi(\bm{W},\varepsilon_{i});w\big)=k\Big\},$
@@ 5.1. Main Settings
Dataset and Model Selections. We evaluate our method on two standard text classification datasets, including AG’s News (AG) (Zhang et al., 2015) and SST-2 (Socher et al., 2013), and further validate generalization across three PLMs: BERT-base-uncased (Devlin et al., 2019), GPT-2 base (Radford et al., 2019), and OPT-1.3B (Zhang et al., 2022).
@@
Two side-by-side plots. Left: effect of the number of watermarked models. Right: effect of the proportion of filtered outliers.
@@
Two side-by-side plots. Left: resistance to model fine-tuning. Right: resistance to model pruning.
@@
Two heatmaps for SST-2. The left heatmap shows verification success rate (VSR) and the right heatmap shows watermark certification accuracy (WCA), comparing different model architectures where rows correspond to benign models and columns correspond to watermarked models.
@@
Figure 10. WSR of our method’s watermarked samples under perturbations in the embedding space. $d_{N}$ and $d_{A}$ denote the noise and adversary-crafted directions, respectively, and $\bullet$’ marks the original watermarked sample.
Two subfigures comparing watermark robustness (WR) of our method under embedding-space perturbations on SST-2. The left panel uses the BadWord trigger and the right panel uses AddSent; the bullet marks the original watermarked sample and the curves show WR under noise and adversary-crafted directions.
@@ 6. Conclusion
In this paper, we revisited existing dataset ownership verification (DOV) methods for pre-trained language models (PLMs) and revealed their vulnerabilities under natural noise and deliberately adversary-crafted perturbations. To address these limitations, we proposed DSSmoothing, a certified dataset watermarking method for textual data based on dual-space smoothing. By introducing controlled perturbations in both the embedding and permutation spaces, DSSmoothing simultaneously enhanced semantic and sequential robustness. Theoretically, we proved that under bounded perturbations, a strict lower bound exists between watermark robustness (WR) and principal probability (PP), enabling provable ownership verification. Experimental results demonstrated that DSSmoothing consistently outperforms existing methods on multiple benchmark datasets and exhibited robustness against potential adaptive attacks, providing a practical and theoretically grounded solution for copyright protection of large-scale text datasets.
Acknowledgements.
This work is supported by Beijing Natural Science Foundation (L251061) and the Industrial Foundation Reengineering and High-Quality Manufacturing Development project (ZC25T320057/100).